How did he end up with the Function at the bottom X2(X1)? I got lost on that last step, i tried simplifying the equation but I could not make it equal the final equation. CONTRACT CURVE : EXAMPLES 1) fiaud (X,,X2), (Y,,Y₂) s.t. MRS*(x..X2)=MRS® 14,42) DATA : E,- E2=10 uA (x1,x2)X, X2 u® (,42) =
Answer #1

x + 4 = 10. Xg ty 2 -10 y = 10-X, 7 = 10 - X2 10-X2 2 (10-x) Zi 20-2x, 20x - 2x, x2 = lux, - X, 42 20x2 – 2x, X2 + x, x9 = 10